OpenStudy (anonymous):

connections academy somebody help me what is 1/7/10 written improper fraction

OpenStudy (anonymous):

Divide!

OpenStudy (anonymous):

but by what 7

OpenStudy (anonymous):

I can't tell what the improper fraction is!

OpenStudy (anonymous):

Can you retype it?

OpenStudy (anonymous):

Use the equation thing.

OpenStudy (anonymous):

what 1/7/10

OpenStudy (anonymous):

is it 17/10

OpenStudy (anonymous):

its 17 over 10

OpenStudy (anonymous):

Do you mean \[\frac{ 1 }{ 7 }\] Do you mean \[\frac{ 7 }{ 10 } or \frac{ 1 }{ 10 } what?\]